OpenStudy (anonymous):

rewrite the following expressions in terms of ln log7^17

jimthompson5910 (jim_thompson5910):

Use the change of base formula \[\Large \log_{7}(17)=\frac{\ln(17)}{\ln(7)}\]

OpenStudy (anonymous):

than alot you're helping me very much tonight